OpenStudy (anonymous):

if f(x)=x-2, for what value of x does f(3x)=f(x+1)

OpenStudy (anonymous):

Well, you have to replace the variable assigned, namely: \[f(3x)=(3x)-2\] and \[f(x+1)= (x+1)-2\], then you equals the equations: and clear x :D

Parth (parthkohli):

\[3x - 2 = (x -2) + 1\]Solve the equality.

OpenStudy (hba):

equate*

OpenStudy (aravindg):

replace x with 3x

OpenStudy (jiteshmeghwal9):

f(x)=x-2 By putting 3x instead of x. we gt f(3x)=3x-2 & by putting x+1 instead of x we gt f(x+1)=x+1-2=x-1 So, f(3x)=f(x+1) => 3x-2=x-1 solve for x :)